Find the solution (intersection point) to the following system:
y=2x-4
y=x+6

Answer by CubeyThePenguin(3113)

y = 2x - 4
y = x + 6

2x - 4 = x + 6
x = 10 ----> y = 2(10) - 4 = 16

intersection: (x, y) = (10, 16)